OpenStudy (anonymous):

The circuit elements have the following values: Vin=18V, R1=1.5Ω, R2=3Ω, R3=4Ω, A=1.5 and B=0.5.

OpenStudy (anonymous):

Q6 The circuit elements have the following values: Vin=18V, R1=1.5Ω, R2=3Ω, R3=4Ω, A=1.5 and B=0.5. Rth???

OpenStudy (nurali):

Question 6 Formula Vout = Vin*(R3-B*R1*R3)/(R3-(B*R1*R3)+R1+R2+A*R2) i1=V0/(R3-B*R1*R3) and Vth is same as Vout. Isc=vIN*(1-B*R1)/(A*R2+R1+R2) Rth=VO/iSC where VO is Vth

OpenStudy (nurali):

If correct please click best response